OpenStudy (anonymous):

Which is a recursive rule for the sequence 9,14,24,44,84? a) a1 = 9, an = a[n-1]+5 b) a1=9, an = a[n-1]+10 c) a1=9, an = 3a[n-1]+13 d) a1=9, an=2a[n-1]-4

OpenStudy (bahrom7893):

So try the first one: a1 = 9 an = a[n-1] + 5, which means: a2 = a[2-1] + 5 = a[1] + 5 = a1 + 5 = 9 + 5 = 14 It fits so far. Repeat for the rest of the numbers.

OpenStudy (anonymous):

Okay got it, would it be A? @bahrom7893

OpenStudy (bahrom7893):

can you post your solution? (The steps you took to check it?)